# Heterogeneous Prompting and Execution Feedback for SWE Issue Test Generation and Selection

### v2

@@ Abstract.
A software engineering issue (SWE issue) is easier to resolve when accompanied by a reproduction test. Unfortunately, most issues do not come with functioning reproduction tests, so this paper explores how to generate them automatically. The primary challenge in this setting is that the code to be tested is either missing or wrong, as evidenced by the existence of the issue in the first place. This has held back test generation for this setting: without the correct code to execute, it is difficult to leverage execution feedback to generate good tests. This paper introduces novel techniques for leveraging execution feedback to get around this problem, implemented in a new reproduction test generator called e-Otter++. Experiments show that e-Otter++ represents a leap ahead in the state-of-the-art for this problem, generating tests with an average fail-to-pass rate of 63% on the TDD-Bench Verified benchmark.
@@
journalyear: 2026copyright: rightsretainedconference: 2026 IEEE/ACM 46th International Conference on Software Engineering; April 15–17, 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Software engineering (SWE) issues are descriptions of a bug or missing feature. SWE issues are usually informal and lack automated tests. But such tests, when they exist, are invaluable for software development. Before issue resolution, tests can reproduce the issue to confirm whether it is real, make it more concrete, and help stakeholders agree upon the desired behavior—following the highly regarded software development practice of test-driven development (TDD) (beck_2002). Once a code change (a patch) is proposed to resolve the issue, running the same reproduction tests helps decide whether to accept the patch. Finally, after issue resolution, adding the tests to the project’s test suite helps prevent future regressions. Reproduction tests for SWE issues should be fail-to-pass ($F\!\!\to\!\!P$): they should fail on the old (pre-patch) code to reproduce the issue and pass on the new (post-patch) code to accept the patch. Unfortunately, generating $F\!\!\to\!\!P$ tests is challenging, because issue descriptions are informal and because the new code patch on which the tests should pass does not yet exist. In addition to helping human software engineers, generating tests from SWE issues can also assist automated SWE agents, as several SWE agents also rely on reproduction tests (arora_et_al_2024; chen_et_al_2024; ehrlich_et_al_2025; jain_et_al_2025; li_et_al_2025; ruan_zhang_roychoudhury_2024; xia_et_al_2025; yang_et_al_2024; zhang_et_al_2024).
@@
There have been several recent attempts at using large language models (LLMs) to automatically generate $F\!\!\to\!\!P$ tests from SWE issues, without access to the resolving code patches. Libro uses an LLM to directly generate 50 tests from Java bug reports, then performs rule-based postprocessing and ranking (kang_yoon_yoo_2023). Other attempts focus on Python, modifying a SWE agent designed to automatically generate code patches to generate tests instead (lin_et_al_2024; mundler_et_al_2024). Recent papers proposed novel LLM-based flows or agents custom-tailored to the problem of generating tests from issues, including AEGIS (wang_et_al_2024), Otter++ (ahmed_et_al_2025), BRT agent (cheng_et_al_2025), and Issue2Test (nashid_et_al_2025). Table 1 shows how e-Otter++ is procedurally distinct. Note that Otter, Issue2Test, and e-Otter do not use inference scaling or test selection.
@@
There are two benchmarks for generating tests from issues: SWT-bench (mundler_et_al_2024) (derived from SWE-bench (jimenez_et_al_2024)) and TDD-Bench Verified (ahmed_et_al_2024-tdd) (derived from SWE-bench Verified (chowdhury_et_al_2024)). SWT-bench Verified, also derived from SWE-bench Verified, is very similar to TDD-Bench Verified. SWT-bench has a leaderboard (https://swtbench.com), topped by the unpublished Amazon Q Developer Agent with an $F\!\!\to\!\!P$ rate of 37.7% on SWT-bench Lite and 49.0% on SWT-bench Verified. However, as there is no paper about Amazon Q Developer Agent, its inner workings are not publicly known. The top system on SWT-bench Verified with a published paper is Otter++ (ahmed_et_al_2025) with an $F\!\!\to\!\!P$ rate of 37.0%, and it also has the exact same $F\!\!\to\!\!P$ rate of 37.0% on TDD-Bench Verified. Given how similar the two verified benchmarks are, this paper evaluates on TDD-Bench Verified only. This paper establishes a new state of the art (SOTA) for SWT-bench Lite with 52.5% $F\!\!\to\!\!P$ rate and for TDD-Bench Verified with 63.0% $F\!\!\to\!\!P$ rate. It does so with a new system, e-Otter++. The “e-” refers to execution feedback: e-Otter++ outperforms prior work by leveraging execution feedback more effectively. Additionally, we apply our approach to the contamination-free benchmark SWE-rebench (badertdinov2025swe) to show the effectiveness of our two primary contributions (execution feedback and heterogeneous prompting), as SWE-bench is suspected to be exposed to recent frontier models during training (liang2025swe).
@@
The first place where e-Otter++ uses execution feedback is after generating a candidate test. The test should fail on the old code, which is easy to accomplish, but it should fail for the right reason, which is challenging. Unlike prior work, e-Otter++ uses execution feedback to retrieve related code and augments the repair prompt with that. The second place where e-Otter++ uses execution feedback is for test selection. After generating and repairing multiple tests, it automatically picks a single final test. The chosen test should pass on the new code, but since the resolving new code is not available, there is no execution feedback to tell whether that is the case. Unlike prior work, e-Otter++ generates candidate code patches as a surrogate for the hidden ground-truth new code. Despite generated code patches being of course imperfect, our experiments show that tests passing on them helps predict passing on the ground-truth code patch. However, even the best test selector is only as good as the pool of generated candidate tests it picks from. Ideally, this test pool should be diverse, and most prior work turns up the LLM decoding temperatures to increase candidate diversity. Unlike prior work, e-Otter++ instead gets diversity from two kinds of systematic prompt variations: issue description morphs, which use an LLM to rewrite the issue text in various ways before test generation, and prompt context masks, which elide various subsets of features from the test-generation prompt. Our experiments show that morphs and masks greatly increase the likelihood of at least one candidate test being $F\!\!\to\!\!P$.
@@ 2. Problem Statement and Background
This section discusses the problem statement and briefly introduces prior work we build upon: Otter (ahmed_et_al_2025), Agentless (xia_et_al_2025), SWT-bench (mundler_et_al_2024), and TDD-Bench Verified (ahmed_et_al_2024-tdd).
@@ 2.2. Reproduction Test Generation: Otter
This paper introduces e-Otter++, which extends an earlier system called Otter++ (ahmed_et_al_2025). Both e-Otter++ and Otter++ generate tests directly from issue descriptions and $c_{\textrm{old}}$. In other words, both are functions genTests that address the problem stated in Section 2.1.
@@ 2.3. Code Patch Generation: Agentless
One of the new techniques in e-Otter++ is how it uses execution feedback from automatically generated candidate code patches. Rather than implementing a new code patch generator, it reuses an existing one, Agentless (xia_et_al_2025), as a component.
@@
Our new system e-Otter++ uses only the first part of Agentless, without the test generation or patch selection of Agentless, as a component (see Section 3.2). While we could have adopted any code patch generator, we chose Agentless because it is open-source and candidate patches are publicly available (agentless_artifact).
@@ 2.4. Test Evaluation Benchmarks
There are two benchmarks for evaluating bug reproduction tests. SWT-bench Lite (mundler_et_al_2024) was derived primarily from the SWE-bench Lite dataset (jimenez_et_al_2024) and contains 276 samples. TDD-Bench Verified (ahmed_et_al_2024-tdd) was derived from SWE-bench Verified (chowdhury_et_al_2024). In TDD-Bench Verified, the authors ran 500 samples from SWE-bench Verified, filtered out samples with zero test coverage or those violating the $F\!\!\to\!\!P$ property of the tests, and ended up with 449 samples. Both benchmarks judge whether a generated test reproduces the described issue by checking whether the test fails before applying the golden patch and passes afterward. In other words, they both implement an evaluation harness consistent with the diagram in Figure 1. One difference between their implementations is that TDD-Bench only runs contributing tests (tests that were added or updated), whereas SWT-bench runs the whole test files that contain the contributing tests. Later, Mündler et al. introduced SWT-bench Verified, which also derives from SWE-bench Verified like TDD-Bench, into their leaderboard (https://swtbench.com). Section 4 reports results on both SWT-bench Lite and TDD-Bench Verified.
@@ 7     $y\leftarrow\textit{LLM}(\textit{repairPrompt}(d_{\textrm{issue}},y,\textit{extraContext})$
As shown in Algorithm 1, the test repair component starts by running an Otter-generated candidate test $y$ on $c_{old}$. It is expected that the test should fail in this phase. However, a test may not fail for the reason mentioned in the issue description. Chen et al. (chen_ma_jiang_2025) reported that Python execution errors during the issue resolution phase correlate with lower resolution rates. They also identified the most prevalent errors—such as ModuleNotFoundError and TypeError—in the generated code, which tend to differ greatly from the reason mentioned in the issue description. To address this, we propose an LLM-based critic that will take the issue description and test log as input and decide whether the test is failing for the right reason. If the critic believes the test is failing for the reason mentioned in the issue description, the repair loop immediately stops. Otherwise, it tries to repair the error with additional information identified by the critic, which includes the possible buggy line, the most relevant text/code snippet from the issue description, and any function that the model wants to see before attempting to write the test again. The critic consists of a single LLM call with temperature 0, where the LLM output contains both the decision and additional information. We explicitly ask for the buggy line because these Python execution errors (e.g., ModuleNotFoundError, TypeError) are localized to one particular line. It is helpful for the model to know which part of the program needs more attention. We have also seen some cases where the issue description has enough hints to avoid execution errors. For example, in “astropy__astropy-14182”, the GPT-4o model generates “ascii.write()”, which is wrong. The correct version is “tbl.write()”, which is not as highly used as “ascii.write()” but is present in the issue description. The critic also proposes function names to read to augment the context (e.g., function arguments, return type), and that actually helps the model in repairing the tests.
After obtaining the additional information from the critic, the repair loop collects the functions from the code that the LLM wants to see. Next, it calls the LLM again to repair the test, feeding the issue description, the current version of the test, the test log, and the additional information collected based on the critic output into the test-repair prompt. Then, the repair loop starts over from this new test as $y$, until the critic is satisfied or for a maximum of 10 iterations. Following prior works (xia_et_al_2025; ehrlich_et_al_2025; chen2021evaluating), the test repair LLM call uses a higher temperature (0.8) to encourage the model to try diverse solutions. We refer to the test generated at this phase as e-Otter test.
@@
Recent works both on generating SWE-patches (ehrlich_et_al_2025; xia_et_al_2025; jain_et_al_2025) and reproduction tests (ahmed_et_al_2025; wang_et_al_2024) have adopted inference scaling. Most works rely on higher temperatures to increase the diversity of solutions. One of the key limitations of such approaches is using the same context for all solutions, which may contain misleading information (e.g., incorrect localization). Also, in a chain of multiple LLM calls, using a higher temperature in each call exponentially increases the number of samples.
Instead of higher temperature, Otter++ (ahmed_et_al_2025) attempted heterogeneous prompting via five masks hiding different parts of contexts and achieved a higher fail-to-pass rate @5. We refer to the five tests from Otter++ masks as planner, full, testLoc, patchLoc, and none. Planner represents the default Otter-generated tests where self-reflective action planning was applied. The other masks partially or completely expose localization information to the model to generate diverse solutions. For example, in ‘testLoc’, the test generation prompt contains only test localization outputs, but no patch localization or planner outputs.
@@
We rerun the Otter pipeline using these new morphed issue descriptions and generate five new sets of tests. Note that issue rewriting may hurt performance since it may remove necessary information and add irrelevant context. However, our goal is to increase the $F\!\!\to\!\!P$ rate @ N, which could be achieved by such rewriting. Prompts and examples for execution-augmented test repair and issue description morphs are in the supplementary material.
@@ 3.2. Code Patch Generator
Our approach uses generated code patches to help with test selection. Prior work does something similar but in the other direction, using generated tests to help with code selection. CodeT (chen_et_al_2023) generates both code samples and corresponding test cases, then executes the generated code samples using the generated test cases and performs a dual execution agreement to choose the best code sample. Agentless (xia_et_al_2025) and later CodeMonkeys (ehrlich_et_al_2025) also use reproduction tests to select patches from multiple candidates but do not evaluate the quality of the reproduction test. Our e-Otter++ uses code patches to evaluate the quality of generated reproduction test candidates and attempt to choose the best one. We use Agentless patches, as opposed to patches from other systems, because their artifacts are open-source and their generated patches are available. Agentless generates a total of 40 code patches (1 greedy at temperature 0 and 39 non-greedy at temperature 0.8) prior to patch selection. We take the normalized version of these 40 patches and select the 5 most frequent ones. We chose 5 considering the number of existing normalized unique patches and our resource constraints.

### v3

@@ Abstract.
A software engineering issue (SWE issue) is easier to resolve when accompanied by a reproduction test. Unfortunately, most issues do not come with functioning reproduction tests, so this paper explores how to generate them automatically. One technique that helps with this task is inference scaling, but traditional temperature-based scaling tends to lack diversity. This paper introduces heterogeneous prompting to address this problem. Another technique that helps with this task is execution feedback, but this is hampered by the fact that the new code (after issue resolution) to execute does not yet exist. This paper introduces novel approaches to get around that problem. We implemented our techniques in a new reproduction test generator called e-Otter++. Experiments show that e-Otter++ represents a leap ahead in the state-of-the-art for this problem, generating tests with an average fail-to-pass rate of 63% on the TDD-Bench Verified benchmark.
@@
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3787837isbn: 979-8-4007-2025-3/2026/04
@@ 1. Introduction
Software engineering (SWE) issues are descriptions of a bug or missing feature. SWE issues are usually informal and lack automated tests. But such tests, when they exist, are invaluable for software development. Before issue resolution, tests can reproduce the issue to confirm whether it is real, make it more concrete, and help stakeholders agree upon the desired behavior—following the highly regarded software development practice of test-driven development (TDD) (Beck, 2002). Once a code change (a patch) is proposed to resolve the issue, running the same reproduction tests helps decide whether to accept the patch. Finally, after issue resolution, adding the tests to the project’s test suite helps prevent future regressions. Reproduction tests for SWE issues should be fail-to-pass ($F\!\!\to\!\!P$): they should fail on the old (pre-patch) code to reproduce the issue and pass on the new (post-patch) code to accept the patch. Unfortunately, generating $F\!\!\to\!\!P$ tests is challenging, because issue descriptions are informal and because the new code patch on which the tests should pass does not yet exist. In addition to helping human software engineers, generating tests from SWE issues can also assist automated SWE agents, as several SWE agents also rely on reproduction tests (Arora et al., 2024; Chen et al., 2024; Ehrlich et al., 2025; Jain et al., 2025b; Li et al., 2025; Ruan et al., 2024; Xia et al., 2025; Yang et al., 2024; Zhang et al., 2024).
@@
There have been several recent attempts at using large language models (LLMs) to automatically generate $F\!\!\to\!\!P$ tests from SWE issues, without access to the resolving code patches. Libro uses an LLM to directly generate 50 tests from Java bug reports, then performs rule-based postprocessing and ranking (Kang et al., 2023). Other attempts focus on Python, modifying a SWE agent designed to automatically generate code patches to generate tests instead (Lin et al., 2024; Mündler et al., 2024). Recent papers proposed novel LLM-based flows or agents custom-tailored to the problem of generating tests from issues, including AEGIS (Wang et al., 2024), Otter++ (Ahmed et al., 2025), BRT agent (Cheng et al., 2025), AssertFlip (Khatib et al., 2025), and Issue2Test (Nashid et al., 2025). Table 1 shows how e-Otter++ is procedurally distinct. Note that Otter, Issue2Test, and e-Otter do not use inference scaling or test selection.
@@
There are two benchmarks for generating tests from issues: SWT-bench (Mündler et al., 2024) (derived from SWE-bench (Jimenez et al., 2024)) and TDD-Bench Verified (Ahmed et al., 2024) (derived from SWE-bench Verified (Chowdhury et al., 2024)). SWT-bench Verified, also derived from SWE-bench Verified, is very similar to TDD-Bench Verified. SWT-bench has a leaderboard (https://swtbench.com), topped (as of when we submitted this paper) by the unpublished Amazon Q Developer Agent with an $F\!\!\to\!\!P$ rate of 37.7% on SWT-bench Lite and 49.0% on SWT-bench Verified. However, as there is no paper about Amazon Q Developer Agent, its inner workings are not publicly known. The top system on SWT-bench Verified with a published paper is Otter++ (Ahmed et al., 2025) with an $F\!\!\to\!\!P$ rate of 37.0%, and it also has the exact same $F\!\!\to\!\!P$ rate of 37.0% on TDD-Bench Verified. Given how similar the two verified benchmarks are, this paper evaluates on TDD-Bench Verified only. This paper establishes a new state of the art (SOTA) for SWT-bench Lite with 52.5% $F\!\!\to\!\!P$ rate and for TDD-Bench Verified with 63.0% $F\!\!\to\!\!P$ rate. It does so with a new system, e-Otter++. The “e-” refers to execution feedback: e-Otter++ outperforms prior work by leveraging execution feedback more effectively. The other core innovation in e-Otter++ is heterogeneous prompting. Additionally, we apply our approach to the contamination-free benchmark SWE-rebench (Badertdinov et al., 2025) to show the generalizability of our innovations, as SWE-bench is suspected to be exposed to recent frontier models during training (Liang et al., 2025a).
@@
The first place where e-Otter++ uses execution feedback is after generating a candidate test. The test should fail on the old code, which is easy to accomplish, but it should fail for the right reason, which is challenging. Unlike prior work, e-Otter++ uses execution feedback to retrieve related code and augments the repair prompt with that. The second place where e-Otter++ uses execution feedback is for test selection. After generating and repairing multiple tests, it automatically picks a single final test. The chosen test should pass on the new code, but since the resolving new code is not available, there is no execution feedback to tell whether that is the case. Unlike prior work, e-Otter++ generates candidate code patches as a surrogate for the hidden ground-truth new code. Despite generated code patches being of course imperfect, our experiments show that tests passing on them helps predict passing on the ground-truth code patch. However, even the best test selector is only as good as the pool of generated candidate tests it picks from. Ideally, this test pool should be diverse, and most prior work turns up the LLM decoding temperatures to increase candidate diversity. Unlike prior work, e-Otter++ instead gets diversity from heterogeneous prompting. It uses two kinds of systematic prompt variations: issue description morphs, which use an LLM to rewrite the issue text in various ways before test generation, and prompt context masks, which elide various subsets of features from the test-generation prompt. Our experiments show that morphs and masks greatly increase the likelihood of at least one candidate test being $F\!\!\to\!\!P$.
@@ 2. Problem Statement and Background
This section discusses the problem statement and briefly introduces prior work we build upon: Otter (Ahmed et al., 2025), Agentless (Xia et al., 2025), SWT-bench (Mündler et al., 2024), and TDD-Bench Verified (Ahmed et al., 2024).
@@ 2.2. Reproduction Test Generation: Otter
This paper introduces e-Otter++, which extends an earlier system called Otter++ (Ahmed et al., 2025). Both e-Otter++ and Otter++ generate tests directly from issue descriptions and $c_{\textrm{old}}$. In other words, both are functions genTests that address the problem stated in Section 2.1.
@@ 2.3. Code Patch Generation: Agentless
One of the new techniques in e-Otter++ is how it uses execution feedback from automatically generated candidate code patches. Rather than implementing a new code patch generator, it reuses an existing one, Agentless (Xia et al., 2025), as a component.
@@
Our new system e-Otter++ uses only the first part of Agentless, without the test generation or patch selection of Agentless, as a component (see Section 3.2). While we could have adopted any code patch generator, we chose Agentless because it is open-source and candidate patches are publicly available (1).
@@ 2.4. Test Evaluation Benchmarks
There are two benchmarks for evaluating bug reproduction tests. SWT-bench Lite (Mündler et al., 2024) was derived primarily from the SWE-bench Lite dataset (Jimenez et al., 2024) and contains 276 samples. TDD-Bench Verified (Ahmed et al., 2024) was derived from SWE-bench Verified (Chowdhury et al., 2024). In TDD-Bench Verified, the authors ran 500 samples from SWE-bench Verified, filtered out samples with zero test coverage or those violating the $F\!\!\to\!\!P$ property of the tests, and ended up with 449 samples. Both benchmarks judge whether a generated test reproduces the described issue by checking whether the test fails before applying the golden patch and passes afterward. In other words, they both implement an evaluation harness consistent with the diagram in Figure 1. One difference between their implementations is that TDD-Bench only runs contributing tests (tests that were added or updated), whereas SWT-bench runs the whole test files that contain the contributing tests. Later, Mündler et al. introduced SWT-bench Verified, which also derives from SWE-bench Verified like TDD-Bench, into their leaderboard (https://swtbench.com). Section 4 reports results on both SWT-bench Lite and TDD-Bench Verified.
@@ 7     $y\leftarrow\textit{LLM}(\textit{repairPrompt}(d_{\textrm{issue}},y,\textit{extraContext})$
As shown in Algorithm 1, the test repair component starts by running an Otter-generated candidate test $y$ on $c_{old}$. It is expected that the test should fail in this phase. However, a test may not fail for the reason mentioned in the issue description. Chen et al. (Chen et al., 2025b) reported that Python execution errors during the issue resolution phase correlate with lower resolution rates. They also identified the most prevalent errors—such as ModuleNotFoundError and TypeError—in the generated code, which tend to differ greatly from the reason mentioned in the issue description. To address this, we propose an LLM-based critic that will take the issue description and test log as input and decide whether the test is failing for the right reason. If the critic believes the test is failing for the reason mentioned in the issue description, the repair loop immediately stops. Otherwise, it tries to repair the error with additional information identified by the critic, which includes the possible buggy line, the most relevant text/code snippet from the issue description, and any function that the model wants to see before attempting to write the test again. The critic consists of a single LLM call with temperature 0, where the LLM output contains both the decision and additional information. We explicitly ask for the buggy line because these Python execution errors (e.g., ModuleNotFoundError, TypeError) are localized to one particular line. It is helpful for the model to know which part of the program needs more attention. We have also seen some cases where the issue description has enough hints to avoid execution errors. For example, in “astropy__astropy-14182”, the GPT-4o model generates “ascii.write()”, which is wrong. The correct version is “tbl.write()”, which is not as highly used as “ascii.write()” but is present in the issue description. The critic also proposes function names to read to augment the context (e.g., function arguments, return type), and that actually helps the model in repairing the tests.
After obtaining the additional information from the critic, the repair loop collects the functions from the code that the LLM wants to see. Next, it calls the LLM again to repair the test, feeding the issue description, the current version of the test, the test log, and the additional information collected based on the critic output into the test-repair prompt. Then, the repair loop starts over from this new test as $y$, until the critic is satisfied or for a maximum of 10 iterations. Following prior works (Xia et al., 2025; Ehrlich et al., 2025; Chen et al., 2021), the test repair LLM call uses a higher temperature (0.8) to encourage the model to try diverse solutions. We refer to the test generated at this phase as e-Otter test.
@@
Recent works both on generating SWE-patches (Ehrlich et al., 2025; Xia et al., 2025; Jain et al., 2025b) and reproduction tests (Ahmed et al., 2025; Wang et al., 2024) have adopted inference scaling. Most works rely on higher temperatures to increase the diversity of solutions. One of the key limitations of such approaches is using the same context for all solutions, which may contain misleading information (e.g., incorrect localization). Also, in a chain of multiple LLM calls, using a higher temperature in each call exponentially increases the number of samples.
Instead of higher temperature, Otter++ (Ahmed et al., 2025) attempted heterogeneous prompting via five masks hiding different parts of contexts and achieved a higher fail-to-pass rate @5. We refer to the five tests from Otter++ masks as planner, full, testLoc, patchLoc, and none. Planner represents the default Otter-generated tests where self-reflective action planning was applied. The other masks partially or completely expose localization information to the model to generate diverse solutions. For example, in ‘testLoc’, the test generation prompt contains only test localization outputs, but no patch localization or planner outputs.
@@
We rerun the Otter pipeline using these new morphed issue descriptions and generate five new sets of tests. Note that issue rewriting may hurt performance since it may remove necessary information and add irrelevant context. However, our goal is to increase the $F\!\!\to\!\!P$ rate @ N, which could be achieved by such rewriting. Prompts and examples for execution-augmented test repair and issue description morphs are in our repository (2).
@@ 3.2. Code Patch Generator
Our approach uses generated code patches to help with test selection. Prior work does something similar but in the other direction, using generated tests to help with code selection. CodeT (Chen et al., 2023) generates both code samples and corresponding test cases, then executes the generated code samples using the generated test cases and performs a dual execution agreement to choose the best code sample. Agentless (Xia et al., 2025) and later CodeMonkeys (Ehrlich et al., 2025) also use reproduction tests to select patches from multiple candidates but do not evaluate the quality of the reproduction test. Our e-Otter++ uses code patches to evaluate the quality of generated reproduction test candidates and attempt to choose the best one. We use Agentless patches, as opposed to patches from other systems, because their artifacts are open-source and their generated patches are available. Agentless generates a total of 40 code patches (1 greedy at temperature 0 and 39 non-greedy at temperature 0.8) prior to patch selection. We take the normalized version of these 40 patches and select the 5 most frequent ones. We chose 5 considering the number of existing normalized unique patches and our resource constraints.